# supplementary

### v1

@@ Abstract.
We investigate spatial confounding in the presence of multivariate disease dependence. In the “analysis model perspective” of spatial confounding, adding a spatially dependent random effect can lead to significant variance inflation of the posterior distribution of the fixed effects. The “data generation perspective” views covariates as stochastic and correlated with an unobserved spatial confounder, leading to inferior statistical inference over multiple realizations. While multiple methods have been proposed for adjusting statistical models to mitigate spatial confounding in estimating regression coefficients, results on interactions between spatial confounding and multivariate dependence are very limited. We contribute to this domain by investigating spatial confounding from the analysis and data generation perspectives in a Bayesian coregionalized areal regression model. We derive novel results that distinguish variance inflation due to spatial confounding from inflation based on multicollinearity between predictors and provide insights into the estimation efficiency of a spatial estimator under a spatially confounded data generation model. We demonstrate favorable performance of spatial analysis compared to a non-spatial model in our simulation experiments even in the presence of spatial confounding and a misspecified spatial structure. In this regard, we align with several other authors in the defense of traditional hierarchical spatial models (Khan and Berrett, 2023; Zimmerman and Ver Hoef, 2022; Gilbert et al., 2025) and extend this defense to multivariate areal models. We analyze county-level data from the US on obesity / diabetes prevalence and diabetes-related cancer mortality, comparing the results with and without spatial random effects.
@@ 1. Introduction
Subsequent articles on spatial confounding have typically adopted a “data analysis” or a “data generation” perspective (Khan and Berrett, 2023). Following Hodges and Reich (2010), spatial confounding arises in the analysis perspective from empirical associations between observed explanatory variables and the eigenvectors of an estimated precision matrix. The proposed adjustments include restricted spatial regression (Hodges and Reich, 2010; Hughes and Haran, 2013), restricted spatial GEE (Hui and Bondell, 2022), and SPOCK Prates et al. (2019). On the other hand, spatial confounding occurs in the generation perspective when covariates $X$ are stochastic, spatially dependent, and correlated with an unobserved confounding effect $Z$ (Paciorek, 2010). The effects on statistical inference are evaluated across the realizations of the observed dataset. Recent methods that address spatial confounding from a data generation perspective include Spatial+ (Dupont et al., 2022), gSEM (Thaden and Kneib, 2018) and a spectral adjustment introduced by Guan et al. (2023).
@@
This manuscript provides new insights into spatial confounding in the context of Bayesian multivariate spatial regression. Recently, Gilbert et al. (2025) proved the consistency of the generalized least squares spatial estimator under a causal inference framework with random sampling locations and a spatial confounding criterion for data generation. We offer an alternative defense of traditional spatial estimation by focusing our attention on statistical inference from both perspectives for finite samples and fixed sampling locations. Furthermore, studies exploring how multivariate dependencies interact with spatial confounding are somewhat limited. Previous work by Urdangarin et al. (2024) analyzes the Spatial+ adjustment method in multivariate settings, while Azevedo et al. (2021) proposes MSPOCK, a multivariate extension that combines the SPOCK framework with a shared component model. We analyze multivariate dependencies in a coregionalized areal model following Jin et al. (2007) and provide theoretical results on the dependence between variables under maximal spatial smoothing. Our results align with several other articles on the defense of traditional hierarchical spatial models (Khan and Berrett, 2023; Zimmerman and Ver Hoef, 2022; Gilbert et al., 2025). We extend the results therein to multivariate areal models.
@@
Subsequently, Section 2 defines the generation and analysis perspectives in a multivariate spatial modeling context. Section 3 compares the effect of adding a spatial random effect on the posterior distribution of fixed effects in the analysis model and provides conditions for the spatial estimator to improve inference under the generation model, given known variance parameters. Section 4 evaluates the analysis models over multiple simulated datasets on a map of Californian counties with a spatial confounder. Section 5 compares non-spatial and spatial analysis of data on obesity and diabetes prevalence at the county level in the US alongside diabetes-related cancer mortality rates. Section 6 concludes with a discussion.
@@ 2.1. Data Generation Model with Spatial Confounder
Let $Y_{i}$ be an $n\times 1$ vector corresponding to the $i$th outcome for each $i=1,\ldots,k$, $X_{1}=[X_{11},\ldots,X_{1p}]$, be an $n\times p$ matrix of predictors, and $Z_{i}$ be an $n\times 1$ vector of unknown confounding effects. We assume data is generated from a model associating $Y_{i}$, $X_{1}$ and $Z_{i}$ as
$Y_{i}=\beta_{0i}1_{n}+X_{1}\beta_{1i}+Z_{i}+\eta a_{i},\quad Z_{i}=\delta_{0i}1_{n}+X_{1}\delta_{1i}+\zeta a_{i},\quad X_{1j}=\mu_{j}1_{n}+\epsilon c_{j},$ (1)
for each $i=1,\ldots,k$ and $j=1,\ldots,p$. The three equations in (1) represent models for $[Y\,|\,X_{1},Z]$, $[Z\,|\,X_{1}]$, and $[X_{1}]$. Scalar intercepts for $Y_{i}$, $Z_{i}$, and $X_{1j}$ are denoted by $\beta_{0i}$, $\delta_{0i}$, and $\mu_{j}$, respectively; $1_{n}$ is the $n\times 1$ vector of ones; and $\beta_{1i}$ and $\delta_{1i}$ are $p\times 1$ vectors of fixed regression coefficients. The distributions in the first two models are specified using $n\times k$ matrices $\eta=[\eta_{1},\ldots,\eta_{k}]$ and $\zeta=[\zeta_{1},\ldots,\zeta_{k}]$ with random zero mean error terms and a shared fixed vector $k\times 1$ $a_{i}$. The distribution in the third model is specified by $\epsilon=[\epsilon_{1},\ldots,\epsilon_{p}]$, which is $n\times p$ with random zero-mean error terms, and constant coefficients comprising the $p\times 1$ vector $c_{j}$. The predictors in $X_{1}$ influence each response $Y_{i}$ directly through $\beta_{1i}$ and through the unobserved confounding effects $Z_{i}$.
[…]
$Y=1_{n}\beta_{0}^{\mathrm{\scriptscriptstyle T}}+X_{1}\beta_{1}+Z+\eta A,\quad Z=1_{n}\delta_{0}^{\mathrm{\scriptscriptstyle T}}+X_{1}\delta_{1}+\zeta A,\quad X_{1}=1_{n}\mu^{\mathrm{\scriptscriptstyle T}}+\epsilon C,$ (2)
where $\beta_{1}$ and $\delta_{1}$ are $p\times k$ with $i$th columns $\beta_{1i}$ and $\delta_{1i}$, respectively; $\beta_{0}=(\beta_{01},\ldots,\beta_{0k})^{\mathrm{\scriptscriptstyle T}}$ and $\delta_{0}=(\delta_{01},\ldots,\delta_{0k})^{\mathrm{\scriptscriptstyle T}}$ are $k\times 1$; and $A=[a_{1},\ldots,a_{k}]$ and $C=[c_{1},\ldots,c_{p}]$ are $k\times k$ and $p\times p$, respectively, and assumed to be invertible. We assume that $\eta$, $\zeta$, and $\epsilon$ are statistically independent and the columns of $\eta$ and $\zeta$ are independent. Hence, $A$ induces the dependence between outcomes in the columns of $Y$ and $Z$, while $C$ specifies the dependence between columns of $X_{1}$. Spatial associations between the columns of $Y$, $Z$, and $X_{1}$ arise from the dependence between the columns of $\eta$, $\zeta$, and $\epsilon$, respectively, or any combination of them.
The vector $\operatorname{vec}(Y)$ is $nk\times 1$ obtained by stacking columns of $Y$ and has variance $\operatorname{Var}(\operatorname{vec}(Y))=(A^{\mathrm{\scriptscriptstyle T}}\otimes I_{n})\operatorname{Var}(\operatorname{vec}(\eta+\zeta))(A\otimes I_{n})+(C^{\mathrm{\scriptscriptstyle T}}\otimes I_{n})\operatorname{Var}(\operatorname{vec}(\epsilon))(C\otimes I_{n})$, where $\otimes$ denotes the Kronecker product and $I_{n}$ is the identity matrix $n\times n$. Since the columns of $\eta$, $\zeta$, and $\epsilon$ are independent, $\operatorname{Var}(\operatorname{vec}(\eta+\zeta))$ is $nk\times nk$ block diagonal with $k$ diagonal block elements, and $\operatorname{Var}(\operatorname{vec}(\epsilon))$ is $np\times np$ block diagonal with $p\times p$ diagonal blocks. We focus on spatially dependent $X_{1}$ and $Z$ by introducing spatial structure in the diagonal blocks of $\operatorname{Var}(\operatorname{vec}(\zeta))$ and $\operatorname{Var}(\operatorname{vec}(\epsilon))$.
For a univariate outcome, the BYM (Besag, York and Mollié) model and the BYM2 reparameterization introduced in Riebler et al. (2016) are flexible classes of spatial analysis models that incorporate spatial and non-spatial error variances. Multivariate areal data analysis for dependent outcomes has a substantial following in epidemiology (see MacNab, 2016, 2018; Gao et al., 2022, and references therein for different approaches). Following Jin et al. (2007), we extend the BYM2 model to the multivariate setting by modeling the error terms of $X_{1}$ and $Z$ as linear transformations of independent latent univariate BYM2 spatial processes. Let $V_{\phi}$ be a fixed positive definite $n\times n$ matrix that encodes spatial autocorrelation, $P$ be a diagonal $k\times k$ matrix with diagonal elements $\rho_{i}\in(0,1)$, and $0_{n}$ be an $n\times 1$ vector of zeroes. The columns of the latent error terms $\{\eta,\zeta,\epsilon\}$ are modeled as
@@
for $i=1,\ldots,k$ and $j=1,\ldots,p$. Then (2) implies $\operatorname{Var}(\operatorname{vec}(Y)\,|\,X_{1})=A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}+A^{\mathrm{\scriptscriptstyle T}}PA\otimes V_{\phi}$. Conditional on $X_{1}$, variation in $Y$ stems from $(\epsilon+\zeta)A$, and $\operatorname{Var}(\epsilon_{i}+\zeta_{i})=(1-\rho_{i})I_{n}+\rho_{i}V_{\phi}$. Thus, $\rho_{i}$ can be interpreted as the spatial proportion of the error variance in the $i$th latent spatial process conditional on $X_{1}$ when $V_{\phi}$ is scaled so that the geometric mean of the marginal variances is approximately one, similar to the univariate BYM2 model (Riebler et al., 2016); $A$ functions both as a scaling and linear transformation from the latent error terms to the observed data. Since $\operatorname{Cov}(Z_{i},\operatorname{vec}(X_{1}))=\delta_{1i}^{\mathrm{\scriptscriptstyle T}}C^{\mathrm{\scriptscriptstyle T}}C\otimes V_{\phi}$, we define $Z_{i}$ as a spatial confounder if $\delta_{1i}\neq 0_{p}$.
@@ 2.2. Spatial and Non-Spatial Bayesian Analysis Models
We estimate $\beta=(\beta_{0},\beta_{1}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ using non-spatial and spatial analysis models. For the former, we employ a conjugate Bayesian multivariate regression analysis model,
$Y=X\beta_{NS}+\nu_{NS},\quad\operatorname{vec}(\nu_{NS})\sim\mbox{N}_{nk}(0_{nk},\Sigma_{NS}\otimes I_{n}),\quad\pi(\beta_{NS},\Sigma_{NS})\propto\mbox{IW}(\Sigma_{NS}\,|\,v,v\Sigma_{0}),$ (4)
where the design matrix $X=(1_{n},X_{1})$ is full column rank, $\beta_{NS}$ is the $(p+1)\times k$ vector of slopes, $\Sigma_{NS}$ is $k\times k$ and encodes between-outcome covariance, and the prior $\pi(\beta_{NS},\Sigma_{NS})$ is uninformative on $\beta$ while placing an inverse-Wishart distribution on $\Sigma_{0}$ with degrees of freedom $v>k-1$ and $k\times k$ scale matrix $\Sigma_{0}$ specifying the mean $\mathbb{E}[\Sigma_{NS}^{-1}]=\Sigma_{0}^{-1}$. We denote parameters in the analysis model with a subscript to emphasize a separation from the fixed unknown parameters in the data generating model in (2). The joint posterior is
$\pi(\beta_{NS},\Sigma_{NS}\,|\,Y,X_{1})=\underbrace{\mbox{IW}(v_{\star},\Sigma_{\star})}_{p(\Sigma_{NS}\,|\,Y,X_{1})}\times\underbrace{\mbox{N}_{pk}\left(\operatorname{vec}\left(\widehat{\beta}_{NS}\right),\Sigma_{NS}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}\right)}_{p(\beta_{NS}\,|\,\Sigma_{NS},Y,X_{1})}\;,$ (5)
where $\widehat{\beta}_{NS}=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Y$, $v_{\star}=v+n$ and $\Sigma_{\star}=v\Sigma_{0}+(Y-X\widehat{\beta}_{NS})^{\mathrm{\scriptscriptstyle T}}(Y-X\widehat{\beta}_{NS})$. Exact samples from $\pi(\beta_{NS},\Sigma_{NS}\,|\,Y,X_{1})$ are obtained by simulating $\Sigma_{NS}\sim\mbox{IW}(v_{\star},\Sigma_{\star})$ and then drawing a value of $\beta_{NS}\sim\mbox{N}_{pk}\left(\operatorname{vec}\left(\widehat{\beta}_{NS}\right),\Sigma_{NS}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}\right)$ for each drawn value of $\Sigma_{NS}$.

### v2

# Spatial Confounding in Multivariate Areal Data Analysis

@@ Abstract.
We investigate spatial confounding in the presence of multivariate disease dependence. In the “analysis model perspective” of spatial confounding, adding a spatially dependent random effect can lead to significant variance inflation of the posterior distribution of the fixed effects. The “data generation perspective” views covariates as stochastic and correlated with an unobserved spatial confounder, leading to inferior statistical inference over multiple realizations. Although multiple methods have been proposed for adjusting statistical models to mitigate spatial confounding in estimating regression coefficients, the results on interactions between spatial confounding and multivariate dependence are very limited. We contribute to this domain by investigating spatial confounding from the analysis and data generation perspectives in a Bayesian coregionalized areal regression model. We derive novel results that distinguish variance inflation due to spatial confounding from inflation based on multicollinearity between predictors and provide insights into the estimation efficiency of a spatial estimator under a spatially confounded data generation model. We demonstrate favorable performance of spatial analysis compared to a non-spatial model in our simulation experiments even in the presence of spatial confounding and a misspecified spatial structure. In this regard, we align with several other authors in the defense of traditional hierarchical spatial models (Khan and Berrett, 2023; Zimmerman and Ver Hoef, 2022; Gilbert et al., 2025) and extend this defense to multivariate areal models. We analyze county-level data from the US on obesity / diabetes prevalence and diabetes-related cancer mortality, comparing the results with and without spatial random effects.
@@ 1. Introduction
Subsequent articles on spatial confounding have typically adopted a “data analysis” or a “data generation” perspective (Khan and Berrett, 2023). Following Hodges and Reich (2010), spatial confounding arises in the analysis perspective from empirical associations between observed explanatory variables and a spatial random effect incorporated into the model to account for spatial variability in the response unexplained by the covariates. The proposed adjustments include restricted spatial regression (Hodges and Reich, 2010; Hughes and Haran, 2013), restricted spatial GEE (Hui and Bondell, 2022), and SPOCK Prates et al. (2019). In the generation perspective, spatial confounding occurs when covariates $X$ are stochastic, spatially dependent, and correlated with an unobserved spatial confounder $Z$ (Paciorek, 2010). Effects on inference are evaluated in realizations of the observed dataset. Recent methods that address spatial confounding from a data generation perspective include Spatial+ (Dupont et al., 2022), gSEM (Thaden and Kneib, 2018) and a spectral adjustment introduced by Guan et al. (2023). Bolin and Wallin (2025) propose covariate smoothing to address generation spatial ”self-confounding” resulting from covariate misspecification.
@@
This manuscript provides new insights into spatial confounding in the context of Bayesian multivariate spatial regression. Recently, Gilbert et al. (2025) proved the consistency of the generalized least squares spatial estimator under a causal inference framework with random sampling locations and a spatial confounding criterion for data generation. We offer an alternative defense of traditional spatial estimation from both perspectives for finite samples and fixed sampling locations. Furthermore, studies exploring how multivariate dependencies interact with spatial confounding are somewhat limited; Urdangarin et al. (2024) analyze the Spatial+ adjustment method in multivariate settings, while Azevedo et al. (2021) propose MSPOCK, a multivariate extension that combines the SPOCK framework with a shared component model. We analyze multivariate dependencies in a coregionalized areal model following Jin et al. (2007) and provide theoretical results on the dependence between variables under maximal spatial smoothing. Our results align with several other articles on the defense of traditional hierarchical spatial models (Khan and Berrett, 2023; Zimmerman and Ver Hoef, 2022; Gilbert et al., 2025). We extend the results therein to multivariate areal models.
@@
Subsequently, Section 2 defines the multivariate data generation and Bayesian analysis models. Section 3 examines the effect of adding a spatial random effect on the posterior distribution of fixed effects in the analysis model and estimation properties under the generation model. Section 4 evaluates the analysis models over simulated datasets on a map of Californian counties with a spatial confounder. Section 5 applies the multivariate Bayesian spatial model to estimates of obesity and diabetes prevalence rates at the county level in the US alongside diabetes-related cancer mortality rates. Section 6 concludes with a discussion.
@@ 2. Data Generation and Analysis Models
2.1. Notation and Definitions
We follow earlier work (Zimmerman and Ver Hoef, 2022; Khan and Berrett, 2023; Gilbert et al., 2025) by distinguishing between the true data generation process and the analysis model used to estimate association between $y_{i}$, the $n\times 1$ vector corresponding to the $i$th outcome, and the $p$ covariates, $x_{1},\ldots,x_{p}$, for $i=1,\ldots k$. Separation of the data generation model with fixed, unknown parameters from the Bayesian analysis model permits examination of model misspecification effects on the frequentist properties of the Bayesian spatial and non-spatial estimators. Table 1 and Table 2 list the variables and parameters in the data generation and analysis models, respectively. Matrices are denoted using uppercase English and Greek letters while vectors and scalars are denoted using lowercase letters.
Variable Dimension Random? Description
$Y$ $n\times k$ Yes Outcome matrix
$X$ $n\times(p+1)$ Yes Design matrix with intercept ($X=[1_{n},X_{1}]$)
$X_{1}$ $n\times p$ Yes Design matrix
$B$ $(p+1)\times k$ No Outcome model covariate regression coefficients
$Z$ $n\times k$ Yes Matrix of spatial confounders
$D$ $p\times k$ No Confounder model covariate regression coefficients
$V_{\phi}$ $n\times n$ No Data generation model latent spatial covariance matrix
AA $k\times k$ No Between-outcome dependence matrix factor of $E_{Y},E_{Z}$
$C$ $p\times p$ No Between-outcome dependence matrix factor of $E_{X}$
$\eta_{j}$ $n\times 1$ Yes Outcome model $j$th latent error vector
$E_{Y}$ $n\times k$ Yes Outcome model latent error matrix
$\zeta_{j}$ $n\times 1$ Yes Spatial confounder model $j$th latent error vector
$E_{Z}$ $n\times k$ Yes Spatial confounder model latent error matrix
$\epsilon_{k}$ $n\times 1$ Yes Covariate model $k$th latent error vector
$E_{X}$ $n\times p$ Yes Covariate model latent error matrix
Table 1. Mathematical notation for variables in the data generation model. Parameters are considered as fixed and unknown quantities to be estimated in the analysis model.
In the data generation model, we assume $y_{i}$ is associated with $z_{i}$, a $n\times 1$ vector of unknown confounding effects, and the $n\times p$ covariate matrix $X_{1}=[x_{1},\ldots,x_{p}]$ according to the model,
$\underbrace{y_{i}=\beta_{0i}1_{n}+X_{1}\beta_{1i}+z_{i}+\sum_{j=1}^{k}a_{ji}\eta_{j}}_{p(y_{i}\,|\,X_{1},z_{i})},\;\;\underbrace{z_{i}=\delta_{0i}1_{n}+X_{1}\delta_{1i}+\sum_{j=1}^{k}a_{ji}\zeta_{i}}_{p(z_{i}\,|\,X_{1})},\;\;\underbrace{x_{j}=\mu_{j}1_{n}+\sum_{k=1}^{p}c_{kj}\epsilon_{j}}_{p(X_{1j})},$ (1)
for each $i=1,\ldots,k$ and $j=1,\ldots,p$. Scalar intercepts for $y_{i}$, $z_{i}$, and $x_{j}$ are denoted by $\beta_{0i}$, $\delta_{0i}$, and $\mu_{j}$, respectively; $1_{n}$ is the $n\times 1$ vector of ones; and $\beta_{1i}$ and $\delta_{1i}$ are $p\times 1$ vectors of regression coefficients. The intercepts and coefficients are fixed, unknown quantities;the $n\times 1$ vectors $\eta_{i}$, $\zeta_{i}$, and $\epsilon_{j}$ are random zero mean latent error terms while $a_{ji}$ and $c_{ji}$ are fixed constants. In Section 3, we subscript expectations as $\mathbb{E}_{DG}[\cdot]$ and variances as $\operatorname{Var}_{DG}(\cdot)$ to indicate evaluation under the data generation model.
Variable Dimension Description
$Y$ $n\times k$ Outcome matrix
$X$ $n\times(p+1)$ Design matrix with intercept
$B_{S}$ $(p+1)\times k$ Covariate regression coefficients (spatial analysis model)
$G$ $n\times k$ Spatial random effects
$\Phi$ $n\times k$ Latent spatial factors
$\Psi$ $n\times k$ Latent non-spatial factors
$M$ $k\times k$ Between-outcome matrix factor (spatial analysis model)
$R$ $k\times k$ Diagonal matrix of spatial variance proportions in latent factors
$W_{\phi}$ $n\times n$ Analysis model latent spatial covariance matrix
$E_{S}$ $n\times k$ Residual error matrix (spatial analysis model)
$B_{NS}$ $(p+1)\times k$ Covariate regression coefficients (non-spatial analysis model)
$E_{NS}$ $n\times k$ Residual error matrix (non-spatial analysis model)
$\Sigma_{NS}$ $n\times k$ Between-outcome covariance matrix (non-spatial analysis model)
Table 2. Mathematical notation for variables in Bayesian non-spatial and spatial analysis models.
In Bayesian analysis, the confounders $z_{1},\ldots,z_{k}$ are not observed and hence the direct effect of the covariates $X_{1}$ on $y_{i}$ is obscured. Since the association between $X_{1}$ and $y_{i}$ is of primary interest, only a model for $y_{1},\ldots,y_{k}\,|\,X_{1}]$ is specified.
@@ 2.2. Data Generation Model with Spatial Confounder
For the data generation model in (1), we focus on spatially dependent $X_{1}$ and $Z$ by introducing spatial structure in $\operatorname{Var}(\zeta_{i})$ and $\operatorname{Var}(\epsilon_{i})$. For a univariate outcome, the BYM (Besag, York and Mollié) model and the BYM2 reparameterization introduced in Riebler et al. (2016) are flexible classes of spatial analysis models that incorporate spatial and non-spatial error variances. Following Jin et al. (2007), we extend the BYM2 model to the multivariate setting by modeling the error terms of $z_{i}$ and $x_{j}$ as linear transformations of independent latent univariate BYM2 spatial processes. Let $V_{\phi}$ be a fixed positive definite $n\times n$ matrix that encodes spatial autocorrelation, $\rho_{i}\in[0,1]$, and $0_{n}$ be an $n\times 1$ vector of zeroes. We assume the error terms $\eta_{i},\zeta_{i},\epsilon_{i}$ are generated independently as
@@
for $i=1,\ldots,k$ and $j=1,\ldots,p$. Since variation in $y_{i}$ conditional on $X_{1}$ stems from $\sum_{j=1}^{k}a_{ij}(\epsilon_{i}+\zeta_{i})$ and $\operatorname{Var}(\epsilon_{i}+\zeta_{i})=(1-\rho_{i})I_{n}+\rho_{i}V_{\phi}$, $\rho_{i}$ can be interpreted as the spatial proportion of the error variance in the $i$th latent spatial process conditional on $X_{1}$ when the geometric mean of the marginal prior variances in $V_{\phi}$ is one (Riebler et al., 2016).
Letting $Y=[y_{1},\ldots,y_{k}]$ and $Z=[z_{1},\ldots,z_{k}]$ be $n\times k$ matrices, we cast (1) as
$Y=1_{n}\beta_{0}^{\mathrm{\scriptscriptstyle T}}+X_{1}B_{1}+Z+E_{Y}A,\quad Z=1_{n}\delta_{0}^{\mathrm{\scriptscriptstyle T}}+X_{1}D_{1}+E_{Z}A,\quad X_{1}=1_{n}\mu^{\mathrm{\scriptscriptstyle T}}+E_{X}C,$ (3)
where $B_{1}$ and $D_{1}$ are $p\times k$ with $i$th columns $\beta_{1i}$ and $\delta_{1i}$, respectively; $\beta_{0}=(\beta_{01},\ldots,\beta_{0k})^{\mathrm{\scriptscriptstyle T}}$ and $\delta_{0}=(\delta_{01},\ldots,\delta_{0k})^{\mathrm{\scriptscriptstyle T}}$ are $k\times 1$; and $A=(a_{ji})$ and $C=(c_{kj})$ are $k\times k$ and $p\times p$, respectively, and assumed to be invertible. $A$ and $C$ function as both scaling and linear transformations from the latent error terms to the observed data. The random residual error matrices $E_{Y}=[\eta_{1},\ldots,\eta_{k}]$ and $E_{Z}=[\zeta_{1},\ldots,\zeta_{k}]$ are $n\times k$ while $E_{X}=[\epsilon_{1},\ldots,\epsilon_{p}]$ is $p\times p$. Since $\operatorname{Cov}(z_{i},\operatorname{vec}(X_{1}))=(\delta_{1i}^{\mathrm{\scriptscriptstyle T}}C^{\mathrm{\scriptscriptstyle T}}C)\otimes V_{\phi}$ where $\otimes$ denotes the Kronecker product and $\operatorname{vec}(X_{1})$ is obtained by stacking the columns of $X_{1}$, we define $z_{i}$ as a spatial confounder if $\delta_{1i}\neq 0_{p}$.
@@ 2.3. Spatial and Non-Spatial Bayesian Analysis Models
2.3.1. Non-Spatial Bayesian Analysis
We estimate $B=(\beta_{0},B_{1}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ using a non-spatial conjugate Bayesian multivariate regression model,
$Y=XB_{NS}+E_{NS},\;\;\operatorname{vec}(E_{NS})\sim\mbox{N}_{nk}(0_{nk},\Sigma_{NS}\otimes I_{n}),\;\;\pi(B_{NS},\Sigma_{NS})\propto\mbox{IW}(\Sigma_{NS}\,|\,v\Sigma_{0},v),$ (4)
where the design matrix $X=(1_{n},X_{1})$ is full column rank, $B_{NS}$ is the $(p+1)\times k$ matrix of intercepts and slopes, $\Sigma_{NS}$ is $k\times k$ and encodes between-outcome covariance, and the prior $\pi(B_{NS},\Sigma_{NS})$ is uninformative on $\beta$ while placing an inverse-Wishart distribution on $\Sigma_{NS}$ with degrees of freedom v>k1v>k-1 and $k\times k$ scale matrix $\Sigma_{0}$ specifying the mean $\mathbb{E}[\Sigma_{NS}^{-1}]=\Sigma_{0}^{-1}$. When $\Sigma_{NS}=I_{k}$, the outcomes $y_{1},\ldots,y_{k}$ are modeled as independent. We denote parameters in the analysis model with a subscript to emphasize a separation from the fixed unknown parameters in the data generating model in (3). The joint posterior is
$\pi(B_{NS},\Sigma_{NS}\,|\,Y,X_{1})=\underbrace{\mbox{IW}(v_{\star},\Sigma_{\star})}_{p(\Sigma_{NS}\,|\,Y,X_{1})}\times\underbrace{\mbox{N}_{(p+1)k}\left(\operatorname{vec}\left(\widehat{B}_{NS}\right),\Sigma_{NS}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}\right)}_{p(\operatorname{vec}(B_{NS})\,|\,\Sigma_{NS},Y,X_{1})}\;,$ (5)
where $\widehat{B}_{NS}=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}Y$, $v_{\star}=v+n$ and $\Sigma_{\star}=v\Sigma_{0}+(Y-X\widehat{B}_{NS})^{\mathrm{\scriptscriptstyle T}}(Y-X\widehat{B}_{NS})$. Exact samples from $\pi(B_{NS},\Sigma_{NS}\,|\,Y,X_{1})$ are obtained by simulating $\Sigma_{NS}\sim\mbox{IW}(v_{\star},\Sigma_{\star})$ and then drawing $\operatorname{vec}(B_{NS})\sim\mbox{N}_{(p+1)k}\left(\operatorname{vec}\left(\widehat{B}_{NS}\right),\Sigma_{NS}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}\right)$ for each value of $\Sigma_{NS}$.
2.3.2. Spatial Bayesian Analysis